I am attepting to solve this equation, but am having difficulty.
2x + 15 = (-2x - 3) / x

Answer by stanbon(48568)

2x + 15 = (-2x - 3) / x
----------------
Multiply both sides by "x" to get:
2x^2+15x = -2x-3
2x^2+17x+2=0
x = [-17 +- sqrt(17^2-4*2*2)]/4
x = [-17 +- sqrt(273)]/4
x = [-17 + 16.5227..]/4 = -0.11932...
OR
x = [-17 - 16.5227..]/4 = -8.380677....
============
Cheers,
Stan H.